You have a rectangular garden with dimensions 7 feet by 12 feet. You want to increase the area of the garden to 176 square feet by increasing each dimension by the same number of feet.
1. Write an equation that can be used to find X, if you want the area to be 176 feet.
2.What is the perimeter of the new garden?
(Having trouble figuring out how to answer this question)Thank You

Answer by rothauserc(4718)

We want the area(A) of the rectangular garden to be 176 square feet after increasing each dimension by x feet.
1) A = (x+7)(x+12) = 176
x^2 + 19x + 84 = 176
x^2 + 19x - 92 = 0
factor above equation
(x+23)(x-4) = 0
x = -23, 4
we use x = 4
our new dimensions are 11 and 16
2) the perimeter(P) of a rectangle is 2*l + 2*w where l is length and w is width
P = 2*11 + 2*16 = 22 + 32
P = 54 feet